# Clusterpath Gaussian Graphical Modeling

### v2

@@ B.2 Derivations of the Gradient and Hessian
$|r_{k\ell}|=\begin{cases}\frac{r_{k\ell}^{2}+\varepsilon^{2}}{2\varepsilon}&\text{if $|r_{k\ell}|<\varepsilon$,}\\ |r_{k\ell}|&\text{otherwise}.\end{cases}$
If $\varepsilon$ is chosen sufficiently small (e.g., $\varepsilon=5\cdot 10^{-3}$), this function closely approximates the absolute value function.
@@
$\frac{\partial L_{\text{sparse}}(b_{kk},\mathbf{r}_{k},r_{kk})}{\partial r_{k\ell}}={}$$\begin{cases}\frac{\mathbf{u}_{k}\mathbf{Z}\mathbf{u}_{\ell}^{\top}}{\varepsilon}r_{k\ell}&\text{if $|r_{k\ell}|<\varepsilon$,}\\ \text{sign}(r_{k\ell})&\text{otherwise}.\end{cases}$
@@
$\frac{\partial^{2}L_{\text{sparse}}(b_{kk},\mathbf{r}_{k},r_{kk})}{\partial r_{k\ell}^{2}}={}$$\begin{cases}\frac{\mathbf{u}_{k}\mathbf{Z}\mathbf{u}_{\ell}^{\top}}{\varepsilon}&\text{if $|r_{k\ell}|<\varepsilon$,}\\ 0&\text{otherwise}.\end{cases}$
@@ Appendix C Additional Simulation Results
This appendix provides supplementary simulation results from the experiments discussed in the main paper. Figures 9 and 10 illustrate the results for the chain simulation design with varying numbers of variables and clusters, respectively. Figure 11 shows the results for designs featuring an approximate block structure. The performance of the methods is evaluated based on estimation accuracy (Frobenius norm), clustering quality (number of clusters and ARI), and sparsity recognition (FPR and FNR).
@@ D.1 S&P 100 Stocks
where $H_{tj}$ and $L_{tj}$ are the high and low prices for stock $j$ during trading day $t$ (Parkinson,, 1980) to study the conditional dependency structure of the stocks’ realized ranges.
As a preprocessing step, we first fit the popular heterogeneous autoregressive (HAR) model of Corsi, (2009) to the individual daily realized range series to capture time dependencies. The HAR model captures the temporal dependence in the daily realized range in a very simple yet parsimonious way, namely by explaining it through a weighted average based on the realized range of the preceding day, week and month. After estimating the HAR models, we obtain the $p=101$ standardized residual series and then apply CGGM and TAGL to these to learn the conditional dependency structure among the stocks; our procedure is in line with Wilms & Bien, (2022). Note that estimating a graphical model to these residuals series is useful in the context of a multivariate time series analysis to capture the contemporaneous relationships among the $p=101$ realized ranges.
@@ D.2 OECD Well-Being Indicators
We analyze the OECD well-being indicator dataset that was previously analyzed in Cavicchia et al., (2022), and we acknowledge the authors for providing us with data access. It contains data on $p=11$ variables related to well-being: education, jobs, income, safety, health, environment, civic engagement, accessibility to services, housing, community, and life satisfaction on which $n=36$ countries are given a score ranging from 0 to 10.
Note, however, that we split the sample of countries into two groups since Cavicchia et al., (2022) have found evidence for two groups of countries ($n_{1}=21$ and $n_{2}=14$) with distinct clustering structures. The first group consists of the countries Australia, Austria, Belgium, Canada, Denmark, Finland, France, Germany, Iceland, Ireland, Italy, Japan, Luxembourg, Netherlands, New Zealand, Norway, Spain, Sweden, Switzerland, United Kingdom, United States, the second group consists of the countries Chile, Czech Republic, Estonia, Greece, Hungary, Israel, Korea, Latvia, Lithuania, Mexico, Poland, Portugal, Slovak Republic, Slovenia, Turkey. In our analysis, we omit Lithuania from the second group due to a missing value, hence $n_{1}=21$ and $n_{2}=14$.
@@ D.3 Humor Styles Questionnaire
We analyze data on the humor styles questionnaire (HSQ) developed by Martin et al., (2003). The $p=32$ items of the HSQ, grouped by the humor style being measured, are listed in Table LABEL:tab:HSQ together with their position in the survey. We use responses to the HSQ on a five-point rating scale (anchored by 1 = “never or very rarely true” to 5 = “very often or always true”), which we obtained from https://openpsychometrics.org/_rawdata/. In addition to the aforementioned $p=32$ items, participants were asked at the end of the questionnaire to indicate the accuracy of their responses. We restrict our analysis to participants who reported that their responses are fully accurate, and after removing 9 observations with missing responses, we retain $n=182$ respondents.
@@
Table 2: List of the $p=32$ items of the humor styles questionnaire of Martin et al., (2003), grouped by the humor style being measured and reported together with their position in the survey. Items marked with an asterisk are reverse-keyed, i.e., the responses on the five-point rating scale are reversed prior to analysis.

### v3

@@ B.2 Derivations of the Gradient and Hessian
$|r_{k\ell}|=\begin{cases}\frac{r_{k\ell}^{2}+\varepsilon_{s}^{2}}{2\varepsilon_{s}}&\text{if $|r_{k\ell}|<\varepsilon_{s}$,}\\ |r_{k\ell}|&\text{otherwise}.\end{cases}$
If $\varepsilon_{s}$ is chosen sufficiently small (e.g., $\varepsilon_{s}=5\cdot 10^{-3}$), this function closely approximates the absolute value function.
@@
$\frac{\partial L_{\text{sparse}}(b_{kk},\mathbf{r}_{k},r_{kk})}{\partial r_{k\ell}}={}$ $\begin{cases}\frac{\mathbf{u}_{k}\mathbf{Z}\mathbf{u}_{\ell}^{\top}}{\varepsilon_{s}}r_{k\ell}&\text{if $|r_{k\ell}|<\varepsilon_{s}$,}\\ \text{sign}(r_{k\ell})&\text{otherwise}.\end{cases}$
@@
$\frac{\partial^{2}L_{\text{sparse}}(b_{kk},\mathbf{r}_{k},r_{kk})}{\partial r_{k\ell}^{2}}={}$ $\begin{cases}\frac{\mathbf{u}_{k}\mathbf{Z}\mathbf{u}_{\ell}^{\top}}{\varepsilon_{s}}&\text{if $|r_{k\ell}|<\varepsilon_{s}$,}\\ 0&\text{otherwise}.\end{cases}$
@@ Appendix C Additional Simulation Results
This appendix provides supplementary simulation results from the experiments discussed in the main paper. Figures 9 and 10 illustrate the results for the chain simulation design with varying numbers of variables and clusters, respectively. Figure 11 shows the results for designs featuring an approximate block structure. The performance of the methods is evaluated based on estimation accuracy (Frobenius norm), clustering quality (number of clusters and ARI), and sparsity recognition (FPR and FNR). Figure 12 displays the total computation time over a grid of values for the aggregation parameter (while keeping other tuning parameters fixed) for an increasing number of variables.
@@ D.1 S&P 100 Stocks
where $H_{tj}$ and $L_{tj}$ are the high and low prices for stock $j$ during trading day $t$ \@BBOPcitep\@BAP\@BBN(Parkinson, 1980)\@BBCP to study the conditional dependency structure of the stocks’ realized ranges.
As a preprocessing step, we first fit the popular heterogeneous autoregressive (HAR) model of \@BBOPcite\@BAP\@BBNCorsi (2009)\@BBCP to the individual daily realized range series to capture time dependencies. The HAR model captures the temporal dependence in the daily realized range in a very simple yet parsimonious way, namely by explaining it through a weighted average based on the realized range of the preceding day, week and month. After estimating the HAR models, we obtain the $p=101$ standardized residual series and then apply CGGM and TAGL to these to learn the conditional dependency structure among the stocks; our procedure is in line with \@BBOPcite\@BAP\@BBNWilms and Bien (2022)\@BBCP. Note that estimating a graphical model to these residuals series is useful in the context of a multivariate time series analysis to capture the contemporaneous relationships among the $p=101$ realized ranges.
@@ D.2 OECD Well-Being Indicators
We analyze the OECD well-being indicator dataset that was previously analyzed in \@BBOPcite\@BAP\@BBNCavicchia et al. (2022)\@BBCP, and we acknowledge the authors for providing us with data access. It contains data on $p=11$ variables related to well-being: education, jobs, income, safety, health, environment, civic engagement, accessibility to services, housing, community, and life satisfaction on which $n=36$ countries are given a score ranging from 0 to 10.
Note, however, that we split the sample of countries into two groups since \@BBOPcite\@BAP\@BBNCavicchia et al. (2022)\@BBCP have found evidence for two groups of countries ($n_{1}=21$ and $n_{2}=14$) with distinct clustering structures. The first group consists of the countries Australia, Austria, Belgium, Canada, Denmark, Finland, France, Germany, Iceland, Ireland, Italy, Japan, Luxembourg, Netherlands, New Zealand, Norway, Spain, Sweden, Switzerland, United Kingdom, United States, the second group consists of the countries Chile, Czech Republic, Estonia, Greece, Hungary, Israel, Korea, Latvia, Lithuania, Mexico, Poland, Portugal, Slovak Republic, Slovenia, Turkey. In our analysis, we omit Lithuania from the second group due to a missing value, hence $n_{1}=21$ and $n_{2}=14$.
@@ D.3 Humor Styles Questionnaire
We analyze data on the humor styles questionnaire (HSQ) developed by \@BBOPcitet\@BAP\@BBNMartin et al. (2003)\@BBCP. The $p=32$ items of the HSQ, grouped by the humor style being measured, are listed in Table LABEL:tab:HSQ together with their position in the survey. We use responses to the HSQ on a five-point rating scale (anchored by 1 = “never or very rarely true” to 5 = “very often or always true”), which we obtained from https://openpsychometrics.org/_rawdata/. In addition to the aforementioned $p=32$ items, participants were asked at the end of the questionnaire to indicate the accuracy of their responses. We restrict our analysis to participants who reported that their responses are fully accurate, and after removing 9 observations with missing responses, we retain $n=182$ respondents.
@@
Table 2: List of the $p=32$ items of the humor styles questionnaire of \@BBOPcitet\@BAP\@BBNMartin et al. (2003)\@BBCP, grouped by the humor style being measured and reported together with their position in the survey. Items marked with an asterisk are reverse-keyed, i.e., the responses on the five-point rating scale are reversed prior to analysis.